The vapor pressure, P, of a certain liquid was measured at two temperatures, T. The data is shown in the table.

T (K) P (kPa)
225 3.59
525 7.61

Keep the pressure units in kilopascals.

If you were going to graphically determine the enthalpy of vaporizaton, ΔHvap, for this liquid, what points would you plot? To avoid rounding errors, use three significant figures in the x‑values and four significant
Answer #1

Ln P versus 1 / T data is plotted:

The slope of the straight line formed is:

m = - 295.83

The change of enthalpy of vaporization is calculated:

ΔH vap = - m * R = - (-295.83) * (0.008314) = 2.45 kJ / mol

If you liked the answer, please rate it positively, you would help me a lot, thanks.